OpenStudy (anonymous):

when a light is incident normally is snells law applicable ? as sin i/sinr becomes 0/0 ??

OpenStudy (anonymous):

If light is incident exactly along the normal, then snells law still does apply but with the angles being zero. The worry here is the division by zero, but remember that snells law is really \(n_1\sin\theta_1=n_2\sin\theta_2\), and hence when \(\theta=0\) both sides equal zero, and all is well. Or to put it more accurately, if \(\theta_1=0\) then we know that the left hand side is zero, and hence, the angle \(\theta_2=0\) and hence no refraction. The only reason why we can express it as a ratio for refraction is because often in problems we are assuming that refraction occurs in the first place because the angle is non-zero, and hence we can divide. The proper way to express it is as the equation above and not as the ratio i.e. as \[n_1\sin\theta_1=n_2\sin\theta_2\] and not \[\frac{n_1}{n_2}=\frac{\sin\theta_2}{\sin\theta_1}\]

OpenStudy (anonymous):

ahh fine but dont know why i dont feel convinced :/ does refraction occur when it is noemal or not ??

OpenStudy (anonymous):

Refraction is the bending of light, either towards or away from the normal. If the incident light is parallel with the normal (i.e. \(\theta=0\)) then the light will carry on through unbent (i.e. the angle of refraction is also zero). Snells, law still holds since\[n_1\sin\theta_1=n_2\sin\theta_2\] and thus \[n_1\sin0=n_2\sin\theta_2\] and hence \(\sin\theta_2=0\) since refractive indices are non-zero, meaning only the angle can be zero. So the rays are not bent (so no refraction in that sense), but Snells law is still applicable.

OpenStudy (anonymous):

no refraction so why even aplly snellslaw ? :/

OpenStudy (anonymous):

you might say it's not a law if it only works sometimes... in this case it works even when the angle is zero - much better than something which works say for every angle except 7, 6.3, 5.999, 34.61423, etc. Snells law is a general statement of what happens - it includes the case "nothing happens"

OpenStudy (anonymous):

Just as Stan has said. It is the same as Newtons second law F=ma. No force means no acceleration, but the law still holds.

OpenStudy (anonymous):

thank you :)